OpenStudy (anonymous):

1. Ethan needs to order T-shirts for all the members. Each T-shirt costs $15 and the shipping fee for the order is $9.99. (b) Ethan has $500 allocated in the club’s budget for T-shirt orders. How many T-shirts can Ethan order with this amount? Show your work and explain your answer.

OpenStudy (texaschic101):

15t + 9.99 <= 500 -- subtract 9.99 from both sides 15t <= 500 - 9.99 15t <= 490.01 -- divide both sides by 15 t <= 490.01/15 t <= 32.66....you have to round it down to 32 because there can't be a .66 of a shirt. check.. 15(32) + 9.99 <= 500 480 + 9.99 <= 500 489.99 <= 500 (correct) Ethan can order 32 or less t-shirts